OpenStudy (hovlaus):

f(x) = −x, x ≤ −1 1 − x^2, −1 < x < 1 x − 1, x ≥ 1 . Evaulate each of the following or explain why it does not exist. lim x→−1 f(x), lim x→0 f(x), lim x→1 f(x) .

OpenStudy (sshayer):

\[find~\lim_{x \rightarrow-1 -}f(x)~and ~\lim_{x \rightarrow-1+ }f(x)\] similarly find \[\lim_{x \rightarrow 1-}f(x)~and~\lim_{x \rightarrow 1+}f(x)\] similarly for x=0

OpenStudy (hovlaus):

Yes! Haha, my prof didn't post solutions so I was wondering if I got the answers right. lim of -1 I got DNE. For lim of 0 I got 1. And for lim of 1 I got 0.

OpenStudy (sshayer):

\[\lim_{x \rightarrow -1-}f(x)=\lim_{x \rightarrow -1-}\left( -x \right)\] put x=1-h,h>0 \[h \rightarrow0~as~x \rightarrow-1-\] \[\lim_{x \rightarrow -1-}f(x)=\lim_{h \rightarrow 0}\left\{ -\left( -1-h \right) \right\}=\lim_{h \rightarrow 0}\left( 1+h \right)=1\] \[\lim_{x \rightarrow -1+}f(x)=\lim_{x \rightarrow -1+}\left( 1-x^2 \right)=1-1=0\] \[\lim_{x \rightarrow -1-}f(x) \neq \lim_{x \rightarrow -1+}f(x)\] \[Hence \lim_{x \rightarrow -1}f(x) ~does~\not~ exist.\] so you are correct. similarly your other limits are also correct.